### v1

# Global Streams, Local Currents: A Data Analysis on Global VOD Content Consumption

Abstract
This study explores global VOD (video on demand) content consumption patterns through a network-based approach. We used Netflix’s ’TV-shows’ ranking data, spanning 822 days across 71 countries, to construct a network where countries are represented as nodes and consumption similarities are reflected as link weights. By applying the Louvain algorithm, we identified three distinct consumption groups: ‘North America and Pan-Europe’, ‘Asia and Middle East’, and ‘Central and South America group’. These groups align closely with geographic, historical, and linguistic divisions, despite no predefined grouping criteria. Notably, Turkiye, often considered a cultural and regional crossroads, exhibited some classification ambiguity but was ultimately grouped with Asia and Middle East. Our findings also reveal the dominant influence of the United States across all groups, while South Korean content, particularly after the success of “Squid Game” in 2021, has gained and sustained popularity in Asia and Middle East and Latin America. This study, based on data, demonstrates that deep-seated cultural histories continue to shape global consumption patterns, even amidst rapid changes in media platforms and content production dynamics.
1 Introduction
Despite of the de-globalization trends in realm of politics and economy, culture is in an unprecedented state of globalization. VOD (Video on Demand) services such as Netflix, are the primary drivers of transnational popular culture. As seen in traditional global media consumption patterns based on cultural proximity[31] however, the consumption patterns of VOD services may heavily depend on geographic proximity. Indeed, Netflix has managed their market by dividing it into four geographic cultural clusters (Europe-Middle East-Africa, Asia-Pacific, The United States-Canada, Latin America).
Despite interest in VOD acceptance[32], empirical research became possible only after Netflix’s TOP 10 consumption data was made public in 2020. Lotz[22] has analyzed the frequency of overlap in the preferred Netflix Top 10 titles by major countries and argued that global consumption of Netflix consists of large consistent markets such as English-speaking and Spanish-speaking regions, and small variable markets such as geographically close Korea, Japan, Taiwan, and the unique market of India. Jang et al.[15] have identified eight groups that consume Netflix similarly based on Netflix’s Top 10 consumption data, discovering that they are differentiated by geographical distance, linguistic similarity, and cultural differences.
The concept of geo-cultural or geo-linguistic markets shows that global VOD services are consumed differently regionally, geographically, and linguistically. According to their diversity report[26] in 2021, Netflix places a strong emphasis on producing content that reflects diverse regions and cultures, in addition to being inclusive in terms of race and gender. The strategy is described as a deliberate response to engage with and cater to unique audiences in different regions[6]. While the geo-cultural and geo-linguistic market concepts do demonstrate that global VOD services are consumed differently regionally, geographically, and linguistically, there remain questions. These include inquiries into how global VOD consumption is defined and separated, which nations are pivotal in this segmentation, and beyond it. It is also essential to persistently monitor the evolving patterns of consumption among similar viewer groups over time with longitudinal data. What distinct viewer groups emerge from the complex interactions of people connected to VOD platforms worldwide?
This study aims to address the question by exploring how global VOD is segmented into different national and regional markets. These divisions may not just by country borders, but also by larger cultural and linguistic areas which can transcend national boundaries. Essentially, we examine how the consumption of VOD services is influenced by and caters to these diverse geo-cultural segments. To give a preliminary report, we found that global VOD content is being consumed in three distinct geo-linguistic consumer groups worldwide. Interestingly, these groups follow the trajectory of human history from the West to the East, from settled to migratory communities (alternatively, from colonizers to the colonized). Some countries are located on the boundary of these consumption groups, with Turkiye being a prime example. Due to its unique historical and geographical location straddling the East and West, Turkiye exemplifies a nation that could fall into more than one of these consumption categories. These patterns are not only geographical but also historical, reflecting human movement and cultural development from ancient to modern times.
This suggests that the ways in which global audiences choose and consume VOD content may be influenced by deep-seated historical and cultural trends. The consumption trajectory of global VOD services, which audiences themselves choose according to their preferences, represents ‘the long future’ predetermined by lengthy human history. ‘The long future’ calls to mind “The Long Revolution”, a monumental work by Raymond Williams[33], author of “Television: Technology and Cultural Form” in 1974[27]. Williams elucidates that unlike the Industrial or the Civil Revolutions, the cultural revolution was a lengthy journey that culminated with the emergence of mass media in the 20th century, long after the spread of literacy and the emergence of a reading public. The term “long future” is used to imply that the historical evolution of human society, including such milestones as the cultural revolution Williams discusses, continues to exert a significant influence on current global trends in VOD media consumption. In other words, the extensive path of our collective past continues to shape how we consume media today, and it suggests that these historical influences will persist as a powerful force shaping media consumption well into the future. Thus, we could call it global streams, local currents. The global VOD consumption patterns shown in this study are not hasty conclusions drawn from a few pieces of data, but rather optimized results discovered through more than a thousand computer simulations of actual VOD consumption data.
2 Materials and methods
This research analyzes the top 10 Netflix show lists for each country from July 2020 to September 2022, to identify global trends and popularity patterns of Netflix Show (NS) and to find similarities in Netflix Show Consumption (NSC) patterns across countries. By examining the differences in content rankings between countries, the ‘NSC distance’ between two countries is calculated, where a shorter distance signifies a closer link between the two countries. We constructed a content consumption network with countries as nodes and classified the countries into groups based on their content consumption patterns using the Louvain algorithm[9]. The Louvain algorithm detects communities within a network by identifying groups with high intra-group edge weights and low inter-group edge weights, determining the optimal number of groups automatically instead of requiring user-specified group numbers. We then determine the countries of origin for popular content within each group, identify the five nations most prolific in producing popular content for each respective group, and further quantify the degree of influence these leading five countries exert on the groups. Finally, by measuring the popularity of content from these countries on a quarterly basis throughout the period, we examine the temporal trends and variations in the consumption share attributed to the top content-producing countries. Below, we provide a detailed account of the data utilized and the methodologies applied in conducting this analysis.
2.1 Data
We analyzed content lists of NS rankings, spanning from the 1st to the 10th place, collected from Flixpatrol over a period of 822 days ($N_{day}$) [^1] July 1, 2020 to September 30, 2022 and across 71 countries ($N_{coun}$) [^2] The 71 countries mentioned are the sum of the members of the three groups listed in footnote 4.. If each content appeared in the list only once across the period, the total number of content entries would be $822\times 71\times 10=583,620$. However, since most content appears in the list across multiple days and in different countries, the actual number of unique content entries was reduced to $N_{cot}=1769$.
We denote the ranking of content $c$ on day $d$ in country $A$ as $g_{A}^{d,c}$. Then country $A$ has a ranking dataset,
$G_{A}=\{g_{A}^{1,1},g_{A}^{1,2},\ldots,g_{A}^{2,1},\ldots,g_{A}^{822,1769}\}$ (1)
composed of $822\times 1769=1,454,118$ data points. From this ranking data $g_{A}^{d,c}$ we introduce score data $s_{A}^{d,c}$ as follows.
$s_{A}^{d,c}=\begin{cases}11-g_{A}^{d,c},&\text{for }g_{A}^{d,c}\leq 10\\ 0,&\text{otherwise}\end{cases}$ (2)
As equation (2) indicates, the 1st place ranking is converted to 10 points, the 10th place to 1 point, and rankings beyond the 10th place are assigned 0 points. By converting ranking data into score data, we can now represent each country as a point in a high-dimensional space of $822\times 1769=1,454,118$ dimensions. The coordinates of a country in this space are given by its score data. For example, the position of country A in this hyperspace is represented by the set of scores, $S_{A}=\{s_{A}^{1,1},s_{A}^{1,2},\ldots,s_{A}^{2,1},\ldots,s_{A}^{822,1769}\}$. Within this coordinate system, we define the distance between two countries as the Euclidean distance in hyperspace, as detailed below.
2.2 NSC distance between countries
We aim to define the ‘NSC distance’ between two countries as the difference in their content consumption patterns throughout the entire analysis period. However, due to heuristic reasons, we first define the daily distance ($dd_{AB,d}$) between two countries, A and B on day d based on the differences in the ranked contents between two countries as follows,
$dd_{AB,d}=\sqrt{\sum_{c=1}^{N_{cot}}{\left(s_{A}^{d,c}-s_{B}^{d,c}\right)^{2}}}$ (3)
In the summation, the terms $\left(s_{A}^{d,c}-s_{B}^{d,c}\right)^{2}$ calculate the squared difference in scores between countries A and B for each piece of content c on day d. This approach ensures that all differences are positive and gives more weight to larger discrepancies, regardless of which country had the higher score. Previous research[15] has focused on whether contents were in the top 10 list or not, without considering their specific rankings. We argue that taking ranking scores into account is crucial because the disparity between the 1st and 10th ranks is often more significant than between listed and unlisted content. This is particularly true given the power-law distribution observed in the popularity of cultural contents[3, 13, 25] which suggests that higher-ranked content has disproportionately greater popularity and influence.
In equation (3), the summation index goes from 1 to $N_{cot}=1769$, but in practice, we only add up at most 20 terms that are not zero. For example, if content c is not in the top 10 for country A, then the corresponding term $s_{A}^{d,c}$ is zero, and the same rule applies to country B. This implies that the only terms you need to consider are those where the content is c is in the top 10 for country A or B. All other terms will be zero and can be ignored in the sum.
The NCS distance ($D_{A,B}$) between two countries, A and B, is defined as the square root of the sum of the squares of daily distances $dd{A}{B}{d}$ between them over the days in the entire analysis period in accordance with the Pythagorean theorem:
$D_{A,B}=\sqrt{\sum_{d=1}^{N_{day}}dd_{AB,d}^{2}}=\sqrt{\sum_{d=1}^{N_{day}}\sum_{c=1}^{N_{cot}}\left(s_{A}^{d,c}-s_{B}^{d,c}\right)^{2}}$ (4)
Some may argue that this distance may not accurately capture the differences in actual NSC patterns between two countries due to the limitations set by equation (4). Specifically, one concern arises because the calculation considers only content ranked in the top 10, excluding any data on ranks below the 10th, which could potentially distort the measurement of consumption differences. However, we believe this limitation is negligible due to the power-law distribution observed in content popularity, which suggests that the most significant insights come from the top rankings. Due to the lack of data for ranks below the 10th, we cannot directly compare the distances calculated using all ranks to those calculated with only the top 10 content. Instead, we assess the robustness of our approach by calculating distances $D_{A,B}$ using only the top 5 ranked contents and comparing this order with the original distances $D_{A,B}$. If these two sets of distances maintain a linear relationship, it confirms that analyzing the top 5 contents is sufficient, thereby supporting the validity of focusing on the top 10 list [^3] For detailed analyses, please refer to the supplementary material.. Another concern arises from the fact that the release dates of content can vary between countries, and in some cases, certain content may be available in one country but not in the other. To verify the validity of our methodology, we recalculated distances using only content that appeared in the top 10 on the same date in both countries. This approach also yielded a strong positive correlation with the distances calculated using our original method, as detailed in the supplementary information. Therefore, we are confident in the validity of our analysis results despite the limitations of approximating countries as points in a hyper-dimensional space and using Euclidean distance to represent differences in VOD consumption across countries.
2.3 NSC Network and Grouping
Research on analyzing digital consumption data, such as that from Netflix, through network-based approaches has been extensively studied in recent years.[4, 20, 1, 21, 12] In our study, we construct a consumption network where the nodes represent countries and the links between them are weighted by the similarity in content consumption between those countries. Given the distances $D_{A,B}$ between each pair of countries, A and B, we define the weights $W_{A,B}$ as follows:
$W_{A,B}=\frac{D_{\text{min}}-D_{A,B}}{D_{\text{max}}-D_{\text{min}}}$ (5)
where $D_{\text{max}}$ and $D_{\text{min}}$ are the maximum and minimum values of $D_{i,j}$ for all pairs of countries $(i,j)$ with $i\neq j$. Through this formula, the link between the two countries that are furthest apart in the hyperspace is assigned a weight of 0, indicating the lowest similarity, while the link between the closest countries receives a weight of 1, indicating the highest similarity.
With this network, we classify the countries into NSC groups using the Louvain algorithm,[9] a method widely applied across various fields for community detection.[8, 23, 18, 7, 11, 16] The Louvain algorithm detects communities by maximizing the modularity $Q$, which is defined as
$Q=\frac{1}{2m}\sum_{AB}\left[W_{A,B}-\frac{k_{A}k_{B}}{2m}\right]\delta(c_{A},c_{B})$ (6)
Here, m represents the sum of the weights of all links, $k_{A}$ and $k_{B}$ are the sums of the weights of the links connected to countries A and B respectively, $c_{A}$ and $c_{B}$ denote the groups to which countries A and B belong, and $\delta(c_{A},c_{B})$ takes a value of 1 if the two countries belong to the same group, and 0 if they belong to different groups. Due to the stochastic nature of the Louvain algorithm in determining the sequence for evaluating nodes for possible reallocation to different groups, we conducted 1,000 simulations to mitigate this variability. Based on these simulations, we assigned each country to the group to which it most commonly was affiliated, thus establishing groups of countries with similar VOD consumption patterns.
3 Results
3.1 VOD similar consumption groups
Figure 1: VOD Similar Consumption Groups. The North America and Pan-Europe group (n = 35) is color-coded in red, the Asia and Middle East group (n = 18) is color-coded in green, and the Central and South America group (n = 18) is color-coded in blue.
Figure 1 shows the countries grouped by the Louvain algorithm from the data of 1,769 pieces of content, indicated by colors. In all 1,000 simulations, the countries of 71 nations were classified into three groups, and the countries belonging to each group were always consistent except for one country. Looking at the countries in each group, it can be confirmed that they are classified into three VOD consumption groups [^4] North America and Pan-Europe Group: Australia, Austria, Belgium, Bulgaria, Canada, Croatia, Cyprus, Czech Republic, Denmark, Estonia, Finland, France, Germany, Greece, Hungary, Iceland, Ireland, Israel, Italy, Latvia, Lithuania, Netherlands, New Zealand, Norway, Poland, Portugal, Romania, Slovakia, Slovenia, South Africa, Sweden, Switzerland, Ukraine, United Kingdom, United State
Asia and Middle East Group: Egypt, Hong Kong, India, Japan, Kuwait, Malaysia, Morocco, Nigeria, Oman, Philippines, Qatar, Saudi Arabia, Singapore, South Korea, Taiwan, Thailand, Turkiye, Vietnam
Central and South America group: Argentina, Bolivia, Brazil, Chile, Colombia, Costa Rica, Dominican Republic, Ecuador, Guatemala, Honduras, Mexico, Nicaragua, Panama, Paraguay, Peru, Spain, Uruguay, Venezuela.
: the ‘North America and Pan-Europe Group,’ the ‘Asia and Middle East Group’, and the ‘Central and South America Group’.
The only exception country was Turkiye. Out of 1,000 instances, Turkiye was classified into the Asia and Middle East group 749 times, and into the North America and Pan-Europe group 251 times. It seems to reflect the national character that is geographically on the border of Europe and Asia and has a strong integrative character culturally. In fact, Turkish TV exerts influence in a variety of markets from the Middle East to Europe and Central Asia[17]. It can be inferred that global VOD consumption is strongly influenced by geographical, linguistic, and historical factors.
Figure 2: Top 5 content producing countries and their scores within each group Note. The background color of the graph represents the group being averaged, and the color of the bars represents the group to which the production country belongs. ‘US’ stands for the Unites States, ‘KR’ for South Korea, ‘CO’ for Colombia, ‘UK’ for the United Kingdom, ‘ES’ for Spain, ‘FR’ for France, ‘JP’ for Japan, ‘TR’ for Turkiye, and ‘MX’ for Mexico.
The collected data comprised a total of 1,769 pieces of content, which were produced by 49 countries. To understand how much content produced by individual producing countries was consumed by each country, we considered the sum of scores of the contents produced by a specific producing country as the score for that producing country. In other words, the score for the United States was the sum of scores of 528 pieces of content, and the score for South Korea was the sum of scores of 203 pieces of content. Since each country has contents ranking from 1st to 10th place daily, converting these to scores and summing them up gives 55 points. That is, if all content from 1st to 10th place in a particular country on a particular date was produced by the United States, then the score for the United States in that country on that date would be 55 points. We calculated the average scores for 71 countries over 822 days. The top 5 producing countries with the highest average scores are shown in Fig. 2-(a). Figure 2-(b),(c) and (d) show the results of calculating the average scores for each of three VOD similar consumption groups. Figure 2-(a) represents the average scores over 822 days for all 71 countries, Fig. 2-(b) for the 35 countries in the North America and Pan-Europe group, Fig. 2-(c) for the 18 countries in the Asia and Middle East group, and Fig. 2-(d) for the 18 countries in the Central and South America group. The background color of the graph represents the group averaged, and the color of the bars represents the group the producing country belongs to.
The United States, which ranks first among all countries (see Fig. 2-(a)), occupies high ranks of first, second, and first respectively in the three groups (each in Fig. 2-(b), (c), (d)). South Korea and Spain, which rank second and fifth overall, are also in the Top 5 across all three groups, indicating that content from the United States, South Korea, and Spain consistently receives global approval as a constant variable. In contrast, Colombia and the United Kingdom, which rank third and fourth overall, are localized to the North America and Pan-Europe group and the Central and South America group, respectively, unlike the aforementioned three countries.
Besides the constant variables of the United States, South Korea, and Spain, the countries that rank highly within each group include the United Kingdom and France in the North America and Pan-Europe group, Japan and Turkiye in the Asia and Middle East group, and Colombia and Mexico in the Central and South America group. Interestingly, all of these are countries whose background color and bar color match, that is, they belong to their respective groups. This shows that countries tend to consume a lot of content from countries within the same group. Ultimately, it can be seen that in global VOD consumption, the United States exercises a strong content dominance, South Korea and Spain have mid-range dominance, and a few countries in each consumption group exert a lower level of dominance.
3.2 The role of key content-producing countries in similar consumption grouping
Figure 3: Similar consumption groups by the leading content-producing countries. (a)Similar consumption groups by US-produced content, (b)Similar consumption groups by South Korea-produced content, (c)Similar consumption groups by Columbia-produced content
Then, to what extent do the key content-producing countries influence the grouping into similar consumption clusters? To understand how much influence the significant content-producing countries have on the formation of consumption clusters, we conducted the Louvain algorithm process again, but this time we only included the consumption data from the leading content-producing country within each of the three identified consumption groups. Networks were constructed from 528 pieces of United States content, 203 pieces of South Korean content, and 27 pieces of Colombian content.
When forming similar consumption clusters by the leading contries’ content, all countries, except Turkiye, always belonged to the same group in 1,000 runs of the Louvain algorithm. Similar to the previous results, most countries consistently belonged to the same group in 1,000 runs of the Louvain algorithm, but some countries were classified into several groups. Nevertheless, all countries belonged to a certain group at least 700 times out of 1,000, and the groups that countries mainly belonged to were represented in Fig. 3.
As seen in Fig. 3, we can observe slight changes in the composition of the earlier three similar groups when we form similar consumption groups around specific country-produced content. When focused on US content consumption, the North America and Pan-Europe cluster remained unchanged. However, the countries in the Asian and Middle East group, as well as the Central and South America group, get reorganized. The newly formed groups are the Asia and Central South American group and the Middle East and Africa group. Within this new classification, Spain, which has previously been in the Central and South America group, is now grouped with the North America and Pan-Europe group. Turkiye, which has previously been in the Asian and Middle East group, is classified into the Middle East and Africa group.
For South Korean content, representing Asia and the Middle East, the content consumption patterns and the overall grouping were similar to the original three groups that were previously identified. Yet, notable differences were observed. For example, Argentina, which has been grouped with the Central and South America group when considering a wide range of content, is now classified alongside North American and Pan-European countries when focus is narrowed to South Korean content. Additionally, both Spain and Turkiye are categorized within the North American and Pan-Europe group.
Finally, when the content consumption patterns are analyzed with Colombian content as the focus, we observed two main cluster groups: a Hispanic Latin America and Spain group and other remaining groups. That is, when looking from the perspective of Colombia, the division of groups appears to be heavily influenced by language, specifically whether the countries are Spanish-speaking or not. This reflects the cultural and linguistic influence on media consumption within the VOD market.
In sum, US content is similarly consumed in Asia, Central and South America, as well as the Middle East and Africa. Specifically, Spain is classified within the North America and Pan-Europe consumption group when it comes to US content, while Turkiye is classified within the Middle East and Africa. This implies that audiences in Spain and Turkiye consume US content in ways that align with the consumption pattenrs of North America and Pan-Europe, and the Middle East and Africa, respectively. South Korean content consumption closely resembles the three similar consumption clusters previously identified. However, countries like Argentina, Spain, and Turkiye appear to consume South Korean content from a North American and Pan-European perspective. Finally, Colombian content is predominantly consumed within the Spanish-speaking cultural sphere, including Latin America’s central western regions and Spain in Europe. All other countries are seen as part of a different cultural sphere. Overall, this analysis suggests that major content-producing countries influence global content consumption patterns, which can align or differ based on cultural and linguistic ties. The way content from these key countries is consumed can help to define or redefine cultural spheres and clusters within the VOD space.
3.3 Influencing trend of content-producing countries
Figure 4: Trends in the rankings of US, South Korean, and Colombian content by similar consumption groups. The red triangle represents the scores for American content, the green circle represents South Korean content, and the blue star represents Colombian content. (a)The Asia and Middle East Group, (b)The North America and Pan-Europe Group, (c)The Central and South America Group
Based on the original analysis, which defined three similar consumption groups by considering all content, we further examined how the content rankings of leading countries within the VOD similar consumption groups, such as the US, South Korea, and Coloumbia, have changed over the data period. This is to determine whether the previously observed content performance of these countries is due to random variation or if it represents a significant and sustained influence. Figure 4 represents graphs that illustrate the quarterly ranking scores of content from the United States, South Korea, and Colombia within the three established similar consumption groups.
In the Asia and Middle East group, there is a close competition in the preference for content from the US and South Korea, but after the third quarter of 2021, when ‘Squid Game’ was released, the preference for South Korean content has consistently increased. In the North America and Pan-Europe group, while American content is significantly preferred, there is a spike in interest for South Korean content around the time ‘Squid Game’ was released, but it soon returns to its original level. In the Central and South America group, South Korean content consumption, which was previously minimal, increases following the appearance of ‘Squid Game’ and maintains a certain level thereafter.
As a result, it can be seen that the global response to South Korean content has changed significantly since the global creative effort of VOD intensified, particularly after 2021 with ‘Squid Game’ as a turning point. In Asia, South Korean content surpasses the consumption of American content, and it shows meaningful growth in the Central and South American region. The temporary reaction in the North America and Pan-Europe group suggests that the cultural force of these groups is strong.
4 Summary and Discussion
4.1 Summary
Our findings show that global countries are categorized into three groups with geographic and linguistic connections based on global VOD consumption distance: the ‘North America and Pan-Europe group,’ the ‘Asia and Middle East group,’ and the ‘Central and South America group.’ In these three groups, the ‘North America and Pan-Europe group’ is most influenced by the United States, the ‘Asia and Middle East group’ by South Korea, and the ‘Central and South America group’ by the United States and Colombia.
Looking specifically at the top 5 countries, the United States is the dominant primary driving nation in all three groups, followed by Spain and South Korea exerting a mid-range level of influence, and other countries standing out based on their economic and cultural power within the respective groups. This partially aligns with the fact that Netflix’s original investment in Europe is made discriminately according to the existing cultural power[2]. It allows us to consider the cultural power of the platform-dominant United States, the distinctive geographical and linguistic expansion of Spain, and the status of so-called culturally advanced countries within each similar group. Exceptionally, South Korea is a special case that cannot be explained geographically or linguistically within this explanatory framework, warranting further investigation.
From the results of content consumption patterns over time, we observed the impact of a highly popular South Korean series, ‘Squid Game’. After ‘Squid Game’, the preference for South Korean content consistently increased in the ‘Asia and Middle East group’ and the ‘Central and South America group’, while in the ‘North America and Pan-Europe group’, the preference for South Korean content quickly subsided due to the overwhelming influence of the United States. These findings indicate that the release of culturally significant content such as ‘Squid Game’ has the potential to reshape global content consumption patterns, even in regions with traditionally dominant preferences. South Korean content, riding on the wave of ‘Squid Game’s’ unprecedented international success, has notably shifted its global standing and consumer appeal, surpassing American content in Asia and marking significant inroads in the Central and South American markets. While this surge in popularity may exhibit some level of retraction post-hype in regions like North America and Pan-Europe, the overall trend suggests a solidifying presence of South Korean media on the world stage, challenging established cultural dominions and reflecting the dynamism of global VOD content consumption.
It is noteworthy to mention Turkiye’s unique position in global media consumption. Turkiye occupies the most ambiguous position among the three similar groups. Data suggests that Turkiye consumes American content from the perspective of the Middle East and Africa group, whereas it consumes South Korean content from the perspective of the North America and Pan-Europe group. Such results suggests that Turkiye’s content preference align variably with different cultural groups depending on the content’s origin.
The consumption behaviors of users on unprecedented global VOD content are a direct reflection of the anthropological trajectory of the world. Raymond Williams explored the gradual progression of cultural revolution in his pivotal work “Television: Technology and Cultural Form” published during the height of the mass media era in 1974. Now, precisely fifty years later in 2024, the predominant consumption of VOD content within the extensive historical and cultural trajectory sends a thrill through researchers. The consumption of global VOD content seems to manifest what Williams might call a ‘long future.’ Yet, it remains to be observed whether the recent success of unique cases such as South Korea, with hits like Squid Game and Pachinko, is an accidental ‘outcome’ or will prove to be a sustained ‘factor’ in the future.
4.2 Discussion
The current findings lead us to realize that VOD provides two significant implications that set it apart from traditional television popular culture. First is regarding the geographic landscape of VOD popular culture. Unlike the 20th-century television that defined a common culture within national borders, VOD culture unfolds over geo-linguistic and regional blocks. Unlike traditional broadcasting, which was bounded by national borders, Netflix represents a transnational, transcultural service that transcends multiple countries and unites them on a platform[21]. That is, VOD platforms have created a new type of common culture that goes beyond traditional geographic and linguistic boundaries. Thus, strictly speaking, VOD services could contribute the popular culture of global regional audiences, reflecting their tastes and, more broadly, the universal values shared among them[5, 14]. It appears that the boundaries of VOD content’s regional blocs are established at the intersection of commonality and universality. VOD services pioneer global landscapes of human interaction, media, and technology, extending beyond the realms of finance and commodities.
The global patterns we see in how people use VOD services point to a shift in the way we understand capitalism. It’s no longer just about a competitive market with many channels, as we saw with postmodern capitalism. Now, we’re entering an era of post-postmodern capitalism, where even though people across different regions and languages consume content differently, there’s a trend towards creating a shared global culture[24]. Post-postmodern culture is all about the richness and diversity that come from our connected world. It’s a culture shaped by everyone getting involved, through the interconnected network of our digital age. The way we watch shows and movies on VOD platforms shows that our personal choices are part of a bigger picture, influenced by our shared history and culture. Furthermore, the way VOD services are tailored to fit what each person wants to watch reflects this new cultural landscape. Nowadays, consuming VOD series from other countries has become a familiar landscape. Just as television consistently transcended the boundaries of the household[30], VOD services, now crossing even national borders, have evolved into a universal culture. This culture reflects the aesthetic preferences of individual viewers and fosters cultural connections with anyone worldwide who shares such cultural ties.
The second point is about the new culture shaped by VOD users worldwide. This universal culture is a complex, self-organized force created by the global participating public, reflecting the political economy unique to VOD platforms that capitalizes on viewers as a resource. Instead of being seen merely as targets for advertising, as was common in past media models, VOD users are now valued for their active engagement and the data they generate, such as their viewing habits and subscription payments[19]. This shift has led to the formation of distinct viewer groups or clusters, each with its own patterns of influence and preferences. This idea moves away from the outdated view of audiences simply as commodities[28] and towards recognizing them as resources within the VOD ecosystem. We should now talk about ‘audience resources.’ The global sourcing of finance, along with the diverse and self-directed behaviors of VOD users, is creating new dynamics in media consumption. The intricate processes of VOD consumption and the activity of viewers as resources are two sides of the same coin within this framework.
Central concepts to complexity theory, such as patterns and self-organization, offer significant conceptual and theoretical resources to platform media, where users act as nodes in a network. As networks perform self-organization, the current research findings can offer insights not only into VOD content consumption leadership within clusters but also into exploring genre similarity within them. The leadership exhibited by certain producing or consuming countries indicates a high probability that genres favored by these countries will be universally consumed within the cluster.
From this perspective, a complex systems approach is both distinctive feature and a limitation of this study. The three similar consumption groups identified here are not static; they can change their patterns at any time depending on the activities of global VOD content producers and the consumption behaviors of global viewers. The consumption of specific countries or genres, and the emotions they invoke, emerges through the interactions of individuals participating as nodes in the vast platform network. Therefore, a media complexity approach to global VOD service research is essential. It is also necessary to consider that the tendency for regional bloc formation could be influenced by local opt-out policies, where VOD services locally differentiate content based on region and country. The cultural power of VOD programming as an independent variable should not be underestimated. Nevertheless, while bloc formation can vary, the dynamism of countries with specific historical and cultural backgrounds appears to transcend such policies. Further research in this area is anticipated.
5 Conclusion
In conclusion, our study delves into the transformative realm of global VOD media consumption, revealing patterns that reflect the complex interplay of cultural, geographical, and linguistic factors. We have observed that VOD platforms, transcending traditional media boundaries, are fostering a novel, self-organized global culture, unconfined by national borders. This new era of post-postmodern capitalism heralds the rise of a so called ‘universal culture’ shaped by a networked audience whose tastes and preferences are as varied as they are interconnected[29]. When we refer to cultural universals—patterns, traits, or elements found commonly across different cultures[10]—universal culture signifies those aspects of life that are present across all human cultures, regardless of race, nationality, or region.
The leadership in content production and consumption that countries like the United States and South Korea have demonstrated suggests the emergence of new forms of cultural influence within the global landscape. Notably, the impact of significant cultural phenomena, such as South Korea’s Squid Game, challenges and redefines traditional media consumption patterns.
This study highlights the importance of considering VOD users as dynamic nodes within a vast network, where each node contributes to a collective culture that is continuously evolving. The implications of this evolution are profound, suggesting a shift away from the audience as a mere commodity towards recognizing viewers as active contributors to the media ecosystem.
The findings underscore the need for an ongoing evaluation of how global VOD content production and consumption may continue to reshape the cultural landscape of societies worldwide. As the global media landscape evolves, so too will the cultural narratives that emerge from this interconnected space. The notion of the ‘long future’ reminds us that the shifts we are witnessing are part of a larger historical continuum, carrying the potential to influence the trajectory of global culture for years to come.
6 List of Abbreviations and Symbols
Symbol/Abbreviation Meaning
NS Netflix Show
NSC Netflix Show Consumption
$N_{day}$ Number of data collection days (822 days)
$N_{coun}$ Number of countries in data collection (71 countries)
$N_{cot}$ Number of contents in data collection (1769 contents)
$g_{A}^{d,c}$ The ranking of content $c$ on day $d$ in country $A$
$G_{A}$ The set of $g_{A}^{d,c}$
$s_{A}^{d,c}$ The score of content $c$ on day $d$ in country $A$
$dd_{AB,d}$ Daily distance between country $A$ and $B$ on day $d$
$D_{A,B}$ The NSC distance between country $A$ and $B$
$W_{A,B}$ The weight between country $A$ and $B$
Table 1: List of abbreviations and symbols used in the manuscript.
7 Declarations
  • Ethics approval and consent to participate
    Not applicable
  • Consent for publication
    Not applicable
  • Availability of data and materials
    The datasets used and/or analysed during the current study available from the corresponding author on reasonable request.
  • Competing interests
    The authors declare no competing interests.
  • Funding
    Ministry of Education of the Republic of Korea and the National Research Foundation of Korea (NRF-2022S1A5A2A03051182).
  • Authors’ contributions
    N.L., J.L., and H.C.J. conceived and planned the project. N.L. conducted the computations, simulations, and generated the figures, with input from all authors. All authors contributed to writing the manuscript.
  • Acknowledgements
    This work was supported by the Ministry of Education of the Republic of Korea and the National Research Foundation of Korea (NRF-2022S1A5A2A03051182).
8 Supplementary information
Figure S1: Comparison of the distances between all 2,485 pairs of countries calculated via two different ways. Across all subfigures, the x-axis denotes distances computed by the original method of Eq. (4) in the main text. The y-axis in each subfigure corresponds to distances recalculated with alternative criteria: (a) assigning a notional distance value of 5.5 to rankings beyond the top 10, (b) considering only top 5 content instead of top 10 content, and (c) including only data that ranked within the top 10 for both countries involved.
In this study, we conducted an analysis on the rankings of Netflix shows, focusing on the top 10 in each country from July 2020 to September 2022. As described in the main text, we collected data on 1,769 content items across 822 days for each country. This data was then transformed into scores: 10 points for 1st place, 9 points for 2nd, decreasing to 1 point for 10th place, and 0 points for any content not featured in the top 10 list. We then modeled each country as a point within a hyper-dimensional space of 822×1769=1,454,118 dimensions, with the country’s coordinates determined by its scoring data. We defined the distance between any two countries as their Euclidean distance within this space and calculate the distances for all ((71×70)/2=2485) pairs of countries. This method offers a straightforward and insightful way to analyze the available data. However, the presence of potential data limitations requires reassessing the validity of our analysis. Variations in content release dates across countries and the availability of specific content only in certain countries can be significant factors to affect the validity of our analysis. Moreover, our analysis does not differentiate between content slightly beyond the top 10 rankings and content that is ranked much lower. To address these concerns, we have introduced alternative definitions for calculating the distance between countries. By comparing these distances, our goal is to determine whether the ordering of countries by distance remains consistent across different definitions, thereby verifying the similarity in the ranking order from one definition to another.
We first examined the appropriateness of assigning a uniform coordinate, ’0’, to content not ranked in the top 10 list. If a content does not appear in the top 10 for both countries, its contribution to the distance between these two countries is nullified, as both countries are assigned the ’0’ coordinate for that content axis. On the other hand, for content that is ranked within the top 10, identical coordinates are allocated only if both countries share the same ranking for that content. However, content not ranked in the top 10 inherently shares the ’0’ coordinate, disregarding actual rank differences beyond the top 10.
To address the possibility that unranked content could have different relative rankings, we propose an alternative distance measure. In this method, we assign a constant value of 5.5 to represent the distances between content ranked 11th or lower. This value mirrors the average score within the top 10 rankings, thereby standardizing the distance for any content not appearing in the top 10 list to an average value of 5.5 uniformly irrespective of its presence or absence in the top 10 list of the other country. Yet, the selection of 5.5 as the constant value is somewhat arbitrary but is used to illustrate that incorporating a specific, finite value for ranks beyond the top 10 does not markedly influence our findings.
Figure S1(a) illustrates the correlation between the distances between countries using two distinct definitions: one (represented on the y-axis) involves assigning a constant value of 5.5 to the distances for content ranked 11th and lower, and the original definition (represented on the x-axis). We observe a strong positive correlation between the distances calculated using the two different definitions. Especially, the ranking order of country pairs based on their distances remains almost consistent.
Another concern arises from our method of exclusively using top 10 rankings to calculate distances, uniformly treating ranks below the 10th due to the unavailability of data beyond the 10th rank. This lack of comprehensive data hinders a direct comparison between distances computed from full rankings and those derived solely from the top 10 content. To validate our methodology, we introduced an alternative calculation for distances focusing only on the top 5 rankings, reducing the content entries to 1,314 pieces from the 1,769 used for the top 10 in the main analysis. We then compared these new results with the original distances based on the top 10 rankings. Figure S1(b) displays this comparison. On the y-axis, distances are calculated from content lists up to the 5th rank, with 5 points assigned to 1st place and 1 point to 5th place, while ranks 6th and below are assigned 0 points. The x-axis shows the original distances as outlined in the main text. The strong positive correlation observed between these two sets of distances supports the reliability of analyses that rely on the top 5 rankings, thereby reinforcing the validity of using only the top 10 rankings for our study.
The variability in content release dates across nations, along with the unique availability of specific content in certain countries, may raise questions regarding our analysis. In order to evaluate the resilience of our methodological approach amidst these challenges, we performed a secondary analysis. We performed an analysis with content that was simultaneously present in the top 10 rankings in both countries being compared. Figure S1(c) illustrates the comparison of distances calculated using data that ranked within the top 10 in both countries (depicted on the y-axis) against distances calculated according to the original methodology (depicted on the x-axis). This analysis reveals a pronounced positive correlation between these two sets of distances, indicating that the relative proximity or remoteness between countries according to our metric is consistently preserved.